how do i solve
x = 7 - 10y
10x - 4y = 18
using the substitution method

Answer by Alan3354(69443)

how do i solve
x = 7 - 10y
10x - 4y = 18
using the substitution method
-----------
x is solved for, x = 7-10y
sub for x in the 2nd equation